OpenStudy (anonymous):

What is the integrating factor of the differential equation (12y+7x)dx + (-6x -4(x^2)(y^(-1))) = 0

OpenStudy (anonymous):

\[(12y + 7x)dx + (-6x - 4x ^{2}y ^{-1})dy = 0\]

OpenStudy (yttrium):

I think it will be helpful if you distribute first the dx's and dy's so you can isolate them later.

OpenStudy (anonymous):

you mean multiply dx and dy by each term in the bracket?

OpenStudy (yttrium):

Yeah.

OpenStudy (anonymous):

\[ \mu=\exp\left(\int \frac{\frac{\partial }{\partial y}(12y+7x)-\frac{\partial }{\partial x}(−6x−4x^2y^{−1})}{−6x−4x^2y^{−1}}dx\right) \]

OpenStudy (yttrium):

Ooops. Sorry.

OpenStudy (anonymous):

the real issue im having is that after ive taken the partial derivative of the first bracket with respect to y and the second bracket with respect to x: